Suppose we have a market demand Q = 18 – P and a cost C(Q) 9) = 3Q?.(10 points) Suppose the two firms cannot collude and instead compete in the Cournot Model in the market described in question
Answer #1

Cournot Model

P = 18 – Q

P = 18 – (Q1 + Q2)

Cost function of Firm 1: C1 = 0.5Q21

Cost function of Firm 2: C2 = 0.5Q21

The profits of the duopolists are

Π1 = PQ1 – C1 = [18 – (Q1 + Q2)]Q1 – 0.5Q21

Π1 = 18Q1 – Q21 – Q1Q2 – 0.5Q21

Π1 = 18Q1 – 1.5Q21 – Q1Q2

Π2 = PQ2 – C2 = [18 – (Q1 + Q2)]Q2 – 0.5Q22

Π2 = 18Q2 – Q22 – Q1Q2 – 0.5Q22

Π2 = 18Q2 – 1.5Q22 – Q1Q2

For profit maximization under the Cournot assumption we have

∂Π1/∂Q1 = 0 = 18 – 3Q1 – Q2

∂Π2/∂Q2 = 0 = 18 – 3Q2 – Q1

The reaction functions are

Q1 = 6 – (1/3)Q2

Q2 = 6 – (1/3)Q1

Replacing Q2 into the Q1 reaction function we get

Q1 = 6 – 1/3[6 – (1/3)Q1]

Q1 = 4.5

And

Q2 = 6 – (1/3)Q1

Q2 = 6 – (1/3)(4.5)

Q2 = 4.5

Thus, the total output in the market is

Q = Q1 + Q2 = 4.5 + 4.5 = 9

And the market price

P = 18 – Q

P = 18 – 9

P = 9

Collusive Model

We have the following information

Demand equation: P = 18 – (Q1 + Q2)

Assuming Q = Q1 + Q2

Total Cost (TC) equation for Firm 1: C1 = 0.5Q21

Total Cost (TC) equation for Firm 2: C2 = 0.5Q22

The main aim of the central agency running the cartel is to maximize the total profit of the cartel

ΠT = Π1 + Π2

Where

Π1 = TR1 – TC1 and Π2 = TR2 – TC2

TR = Total Revenue

TC = Total Cost

Thus

ΠT = (TR1 + TR2) – TC1 – TC2

ΠT = P(Q1 + Q2) – 0.5Q21 – 0.5Q22

ΠT = (18 – Q1 – Q2)(Q1 + Q2) – 0.5Q21 – 0.5Q22

ΠT = 18Q1 – Q12 – Q1Q2 + 18Q2 – Q1Q2 – Q22 – 0.5Q21 – 0.5Q22

ΠT = 18Q1 + 18Q2 – 2QAQB – 1.5QB2 – 1.5QA2

Setting the partial derivatives equal to zero we obtain

∂Π1/∂Q1 = 18 – 3Q1 – 2Q2 = 0

∂Π2/∂Q2 = 18 – 3Q2 – 2Q1 = 0

Q1 = 6 – (2/3)Q2

Q2 = 6 – (2/3)Q1

Solving for Q1 and Q2 we obtain

Q1 = 3.6

Q2 = 3.6

Total Output: Q = Q1 + Q2

Q = 3.6 + 3.6 = 7.2

P = 18 – Q

P = 18 – 7.2

P = 10.8

So, we can see that in Collusive model the price is higher and output is lower as compared to the Cournot model. So, we can say that while consumers will prefer Cournot model, the two firms will prefer Collusive model.